# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 1 Introduction
The study of self-gravitating fluids is a classical problem in mathematical physics, closely related to the modeling of stars and gaseous astrophysical bodies. The governing model is the Euler–Poisson system, which has been extensively analyzed under various assumptions and classes of equations of state:
@@
where $\rho(x,t)\geq 0$ is the density of fluids at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ is the gravitational potential. Pressure $P(\rho)$ depends on the density only, with some appropriate assumptions similar to McCann’s [16], which will be described later.
While recent research has extended to rotating binary systems—such as McCann’s binary stars [16] and the star-planet systems explored in the authors’ accompanying work [6], a thorough understanding of the simpler, non-rotating single-body problem remains very important. The properties of these isolated bodies are often inspiring and provide technical tools required to handle the two-body case.
”Non-rotating” implies that the velocity $v$ is zero and the density $\rho$ is time-independent (stationary). This configuration can be regarded as a special case of the uniformly rotating system with the angular momentum is zero. Under these conditions, the Euler-Poisson equations can be reduced to the following equation:
@@
where $\tilde{\rho}$ is a density function (in general with compact support), and $V_{\tilde{\rho}}(x)={\int{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}d^{3}y}}$.
Early rigorous works on the non-rotating single-body problem include Auchmuty and Beals [1], who studied the rotating case but also covered the non-rotating situation by setting the angular momentum to zero. They analyzed variational problems associated with gravitational energy minimization. Lieb and Yau [13] contributed results on uniqueness and other properties of ground states within the framework of quantum mechanics. These results were later synthesized and extended by McCann [16]. However, McCann did not provide detailed proofs, and some of the propositions are not entirely straightforward to justify. In one of the author’s other works [7], the conversion from the Euler–Lagrange equation to the Euler–Poisson equation was rigorously established, with particular attention given to showing that the energy-minimizing solutions still satisfy the Euler–Poisson equations at their boundaries.
@@
The results of this paper are adapted from the author’s Master’s thesis [5], which also gave rise to two companion papers: [7] and [6]. The present work holds intrinsic interest within the theory of self-gravitating fluids. Furthermore, serving as a bridge within the thesis, this work builds upon conclusions from [7] and, in turn, provides essential a priori estimates for the study of rotating star–planet systems in [6], where comparisons to the non-rotating case are essential.
@@ 2 Construction and Results of Non-Rotating Systems
In this section we revisit and adapt the results theorems for non-rotating stars given by Auchmuty and Beals [1] and Lieb and Yau [13], based on McCann’s or Author’s construction and assumptions [16] [7]. In the first subsection, we introduce some basic settings, notations, assumptions and problem settings such as variational formulation. In the second subsection, we review the existence theorems for non-rotating stars given by Auchmuty and Beals [1] and McCann [16]. In the third subsection, we adapt the uniqueness results from the quantum mechanical framework of Lieb and Yau [13] to the classical Newtonian mechanical setting
2.1 Notations and problem setting
We introduce some notations and problem settings based on authors’ another paper [7] (which are also similar to those in McCann [16] and Jang and Seok [11]).
Definition 2.1 (Notations).
We give the following definitions:
  1. (i)
    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
  2. (ii)
    Nonnegative (Positive) part $\lambda]_{+}:=\max\{\lambda,0\}$.
  3. (iii)
    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
  4. (iv)
    $\mu$: Lebesgue measure.
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)=0$ on $\mathbb{R}^{3}$ for non-rotating case. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we follow Auchmuty and Beals’s assumptions [1] and first consider a general form of the pressure $P(\rho)$ as the following:
@@ 2.1 Notations and problem setting
Assume $\sigma$ is an energy minimizer, in order to show $\sigma$ is the solution to (EP’) in classical sense, we want the differentiability of $\sigma$ or of $P(\sigma)$, hence we introduce the following assumptions:
@@
Actually in [7] we introduces another more general but also somewhat technical assumption:
@@
Due to (F2) we know $A(s)$ is well-defined. We also list some properties of $A(s)$ from authors’ another paper [7, Section 2].
Lemma 2.2 (Properties of $A$ [7, Section 2]).
@@
    Moreover, $A^{\prime}(s)s-A(s)=P(s)$, $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}\geq 0$ a.e..
@@
    The inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous.
@@
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use it.
We consider a “potential admissible class” for $\rho$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (4)
@@
$U(\rho):=\int A(\rho(x))d^{3}x$ (7)
$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\rho(x)\sigma(y)}{|x-y|}d^{3}yd^{3}x$ (8)
Here $U(\rho)$ is called internal energy.
@@
$E(\rho)=E_{0}(\rho)+\frac{1}{2}\int\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)d^{3}x$ (9)
@@
In this paper we mainly consider the non-rotating case ($L=0$ or $J=0$, reflected in the subscript 0{0} in $E_{0}(\rho)$). This is not only for the sake of simplicity but also help to produce more results such as uniqueness result in subsection 2.3.
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [7, Proposition 2.10]. Hence $E_{0}(\rho)$ is well-defined.
@@
$mR\left(\mathbb{R}^{3}\right)$ $=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=m\right\}$ (12)
@@
From Remark 2.4, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, selecting the specific exponent $\frac{4}{3}$ can also be shown as a natural assumption if we hope the minimal energy $\inf\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)$ to be finite. We will discuss this later (Proposition 2.17, see also [1, Section 1,6 and 8], and [16, Section 2]). We will also show actually $\frac{4}{3}$ is the critical number such that maximal energy is finite (Corollary 2.19).
@@
Since the energy is translation-invariant, we may just consider looking for a energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, where
$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (14)
@@ 2.2 Existence Results for Non-rotating Bodies
Recall that in the calculus of variations, if an energy functional has minimizer, then this minimizer must satisfy the Euler–Lagrange equation [^1] When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem A.1 or [7, Section 2].. Moreover, it turns out that the Euler–Lagrange equation is equivalent, in a certain sense, to the Euler–Poisson equation (see for example, [7, Section 2] or [11]). Based on this observation, we have Auchmuty, Beals, Lieb and Yau’s results. They not only show the existence of one-body and non-rotating solutions of Euler-Poisson equations, but also indicate some useful properties. McCann lists those results as the following theorem:
@@
    on all of $\mathbb{R}^{3}$ and a single Lagrange multiplier $\lambda_{m}\left\{\begin{array}[]{l}<0,m>0\\ =0,m=0\end{array}\right$.;
@@
Given that $\rho$ is an energy minimizer, to exploit this property better, we introduce the following perturbation set $P_{\infty}(\rho)$(the motivation behind its definition and related properties can be found in [7, Section 4]), which depends on $\rho$:
@@
Lemma 2.8 (Differentiability of Energy $E_{0}(\rho)$ [7, Lemma 5.3, Remark 5.4]).
Given $\rho\in m{R}^{3}$ with $U(\rho)<\infty$, we have $E_{0}(\rho)$ is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. The derivative at $\rho$ is $E_{0}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{0}^{\prime}(\rho)(\sigma)=\int E_{0}^{\prime}(\rho)\sigma$ [^2] To remain consistent with the notation in [1, Section 4], we use $E_{0}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{0}^{\prime}(\rho)$ is given by
@@
To apply Lemma 2.8, we need to verify $U(\rho)<\infty$. It turns out if $\rho\in L^{1}\cap L^{\infty}$, or $\rho$ minimizes $E_{0}(\rho)$ locally (under the topology induced by Wasserstein $L^{\infty}$ distance) or globally on $m{R}\left(\mathbb{R}^{3}\right)$, then $U(\rho)<\infty$, see [7, Section 5].
@@
To prove the existence of minimizer  [^3] It is nature to consider direct method [9]. However, so far, we only know $mR\left(\mathbb{R}^{3}\right)$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, which only implies weak star convergence. It is not enough to show the weak continuity of $G(\rho,\rho)$ and then the weak lower semi-continuity of $E_{0}(\rho)$ with respect to $\rho$ on $W_{\infty}\coloneq m{R}\left(\mathbb{R}^{3}\right)$, see Proposition A.7 or Proposition A.8. Hence, direct method may not be valid on $W_{\infty}$. We need to construct a constrained admissible class $W_{R}$, we first construct a constrained admissible class
@@
Notice that $B_{R}(0)$ is bounded set, by Proposition A.7 or Proposition A.8 we can show the weak continuity of $G(\rho,\rho)$ with respect to $\rho$ on $W_{R}$. Then one can apply direct method to show the existence of minimizer $\sigma_{R}$ (depending on $m$) on $W_{R}$. In order to show such minimizer $\sigma_{R}$ is truly a minimizer on the whole admissible class $W_{\infty}$ for sufficient large $R$, we need the following 4 lemmas.
@@
    Show $\forall R>\widetilde{R}_{0},E_{0}\left(\sigma_{R}\right)\leq a<0,\lambda_{R}\leq l<0$, where $\lambda_{R}$ is the corresponding Lagrange multiplier. (Lemma 2.25)
@@
The proofs of the above four lemmas are postponed to later text. Once these lemmas are established, we obtain a minimizer which satisfies Parts (iv) (v). Therefore, Part (i) holds true with finite $e_{0}(m)$ by the second lemma. With the similar strategies similar to that employed in the proofs of the above lemmas, one can show Part (iv) and Part (v) actually hold true for any minimizer $\sigma_{m}$ (observe that the last term in (26) can be replaced by $\frac{m}{4R}$ since $\int_{B_{R}(0)}\sigma_{m}\geq\frac{m}{2}$ for sufficient large $R$). More precisely, the proof of Part (v) also relies on Part (vi) to show that outside the ball of radius $R_{0}(m)$, $\sigma_{m}$ is not only almost everywhere zero but truly zero everywhere. Moreover, we have the fact that the Lagrange-multiplier $\lambda_{m}$ is negative, which follows from the similar arguments in the proof of Step (1) (Lemma 2.22) showing $\lambda_{R}<0$, see also Remark 2.26.
Now we have the existence of minimizer $\sigma_{m}$. Since $\sigma_{m}$ is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, the set of admissible perturbations here is exactly $P_{\infty}\left(\sigma_{m}\right)$. One can apply (generalized) Lagrange multiplier theorem (Theorem A.1), together with the fact that there are sufficient many functions in $P_{\infty}(\sigma_{m})$, to obtain the following relations (see also [1, Lemma 2] or [7, Subsection 5.2]):
@@
$E_{0}^{\prime}\left(\sigma_{m}\right)=\lambda_{m}\text{ a.e. where }\sigma_{m}(x)>0.$ (19)
@@
$A^{\prime}\left(\sigma_{m}(x)\right)\leq V_{\sigma_{m}}(x)+\lambda_{m}\text{ a.e. where }\sigma_{m}(x)>0$ (20)
Due to Lemma 2.2 (iii), notice $\lambda_{m}<0$, we can obtain $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. As the proof of [7, Theorem 2.25 (iii)], we can apply bootstrap method to show $V_{\sigma_{m}}$ is continuously differentiable. On the other hand (18)(19) implies
@@
where $\phi=\left(A^{\prime}\right)^{-1}$ is continuous. This means $\sigma_{m}$ has a continuous representative. Moreover, if $P(\rho)$ satisfies (F4), then $\phi\in C^{1}((0,\infty))$ by Lemma 2.2 (vii), which implies $\sigma_{m}\in C^{1}\left(\left\{\sigma_{m}>0\right\}\right)$ as we show in the proof of Theorem [7, Theorem 2.25 (vi)]. (Part (vi)).
@@
Part (iii) comes from a strong rearrangement inequality in Lieb [14, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [21]. One can also refer to [16, Proof of Theorem 6.1] or [6, Lemma 3.3.3]. (Part (iii))

### v2

@@ 1 Introduction
The motion of stars and planets has long been a classical subject in celestial mechanics, astrophysics, and mathematics. Historical models such as heliocentrism, proposed by Copernicus, and later studied by Newton and others on stellar rotation, form the foundation of rotating star or planet models. Early developments by Maclaurin, Jacobi, Poincaré, and Liapounov focused on incompressible, homogeneous or nearly homogeneous bodies like the Earth. For gaseous bodies such as Jupiter, stars, or systems like galaxies, density inhomogeneity and compressibility need to be accounted for. In these contexts, such bodies can be modeled as self-gravitating fluids, governed by the Euler–Poisson equations which have been extensively analyzed under various assumptions and classes of equations of state. We consider a fluid with density $\rho(x,t)\geq 0$, velocity $v(x,t)\in\mathbb{R}^{3}$, and gravitational potential $V(x,t)\in\mathbb{R}$, where $x\in\mathbb{R}^{3}$ is position and $t\geq 0$ is time. Assuming the pressure $P(\rho)$ depends on the density only (with appropriate assumptions to be described later), the Euler–Poisson system is given by
@@
The first attempts to construct axisymmetric rotating solutions to the compressible Euler–Poisson equations were made by astrophysicists Milne [47] and Zeipel [59] in the early 1920s, followed by Chandrasekhar [11] and Lichtenstein [41] in 1933. Further historical accounts can be found in [13, 35].
Significant progress has been made for the single rotating star problem, where the star rotates about its axis. A special case of the rotating system is when the velocity $v$ is zero. In this case, the system is actually non-rotating, and the density $\rho$ is time-independent (stationary). Under these conditions, the Euler-Poisson equations reduce to the following equation:
@@
where $\tilde{\rho}$ is a density function (in general with compact support), and $V_{\tilde{\rho}}(x)={\int_{\mathbb{R}^{3}}{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}\,dy}}$.
For the non-rotating case, when the pressure satisfies a polytropic law, the solutions are known as Lane–Emden solutions. However, to obtain solutions for more general pressure laws and more general angular momentum (thus including the rotating case), two approaches have been employed. The first is variational method, as used by Auchmuty and Beals [3, 4] to establish the existence of axisymmetric equilibria for compressible fluids; see also [2, 4, 9, 14, 40]. The second is a perturbative method based on the implicit function theorem, applied near non-rotating stars [31, 32, 30, 55, 56]. Nonlinear dynamical stability of rotating stars was shown in [44, 45] via variational techniques, while stability theory for non-rotating stars can be found in [20, 34, 50].
Each method has its advantages and limitations. For example, the variational approach typically requires $\gamma>\frac{4}{3}$ when $P(\rho)=K\rho^{\gamma}$ to prevent gravitational collapse (see Remark 2.13 and Proposition 2.16 of this paper, [46, Section 2], or [4, Section 1, 6 and 8]). In contrast, perturbative methods can handle the range $\frac{6}{5}<\gamma<2$ [32], but they rely on prior knowledge of solutions in special non-rotating cases.
These methods are quite general and apply not only to single-body problems but also to multi-body systems. Variational methods were used by McCann [46] for binary stars and star-planet systems in the authors’ accompanying work [16]. Perturbative approaches appear in the work of Alonso-Orán, Kepka, and Velázquez [1] on incompressible Euler–Poisson systems with an external particle, as well as in Lichtenstein’s construction of rotating binary stars [41].
Researchers have also considered systems with more than two stellar objects, such as galaxies or general $N$-star systems. For rotating galaxies, single rotating galaxies were constructed in [53, 55] in the spirit of Lichtenstein and Heilig [30, 41]. There is extensive literature on non-rotating galaxies: see [6, 7, 26, 51] and references therein. The orbital stability of stationary solutions has seen considerable progress over the past two decades [28, 23, 24, 25, 22, 27, 36, 37, 38, 39, 51, 58], with recent work on linearly oscillating galaxies in [29]. In [10], Campos, Del Pino, and Dolbeault constructed $N$-body rotating galaxies by connecting them to relative equilibria (point particles) in $N$-body dynamics with small uniform angular velocity. They perturbed radial equilibria, and solved for the gravitational potential rather than the density, using a finite-dimensional reduction. Although natural, Jang and Seok remark in [33] that the framework in [10] is not ideal for stability analysis. They adapt Rein’s reduction method [50, 51] to extend the discussion of binary star solutions to establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [33]. Furthermore, they note that for $N\geq 3$, uniformly rotating $N$-body stellar configurations lack a variational structure analogous to the binary case and are not generally expected to be stable [33].
While recent research has extended to rotating multi-body systems, a thorough understanding of the simpler, non-rotating single-body problem remains very important. In addition to the existence result established via variational methods by Auchmuty and Beals [4], Lieb and Yau [42] contributed results on uniqueness and other properties of ground states within the framework of quantum mechanics. The properties of these isolated bodies are often inspiring and provide technical tools required to handle the two-body case. These properties were hence later synthesized and extended by McCann [46]. However, McCann did not provide detailed proofs, and some of the propositions are not entirely straightforward to justify.
@@
The results of this paper are adapted from the author’s Master’s thesis [15], which also gave rise to two companion papers: [17] and [16]. The present work holds intrinsic interest within the theory of self-gravitating fluids. Furthermore, serving as a bridge within the thesis, this work builds upon conclusions from [17] and, in turn, provides essential a priori estimates for the study of rotating star–planet systems in [16], where comparisons to the non-rotating case are essential. In [17], the conversion from the Euler–Lagrange equation to the Euler–Poisson equation was rigorously established, with particular attention given to showing that the energy-minimizing solutions still satisfy the Euler–Poisson equations at their boundaries. Nevertheless, this work is largely self-contained. Results from [17] are given with proof outlines herein or cited as isolated technical details with complete proofs in [17].
@@ 2 Construction and Results of Non-Rotating Systems
In this section we revisit and adapt the results theorems for non-rotating stars given by Auchmuty and Beals [4] and Lieb and Yau [42], based on McCann’s or the author’s construction and assumptions [46] [17]. In the first subsection, we introduce the problem settings and some basic properties. In the second subsection, we review the existence theorems for non-rotating stars given by Auchmuty and Beals [4] and McCann [46]. In the third subsection, we adapt the uniqueness results from the quantum mechanical framework of Lieb and Yau [42] to the classical Newtonian mechanical setting.
2.1 Problem settings
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)=0$ on $\mathbb{R}^{3}$ for non-rotating case. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we consider a general form of the pressure $P(\rho)$ as the following (which is consistent with [17], see also [4, 46, 33]):
Assumption 2.1.
@@ 2.1 Problem settings
Assuming that $\sigma$ is an energy minimizer, in order to show $\sigma$ is the solution to (EP’) in the classical sense, we want the differentiability of $\sigma$ or of $P(\sigma)$, hence we introduce the following assumption:
@@
Actually in [17] we introduce another more general but also somewhat technical assumption:
@@
Due to (F2) we know $A(s)$ is well-defined. We also list some properties of $A(s)$ from another paper by the author [17, Section 2].
Lemma 2.2 (Properties of $A$ [17, Section 2]).
@@
    Moreover, $A^{\prime}(s)s-A(s)=P(s)$, $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}\geq 0$ a.e.;
@@
    The inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous;
@@
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use them.
We consider an “admissible class” for $\rho$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ (4)
@@
$U(\rho):=\int_{\mathbb{R}^{3}}A(\rho(x))\,dx$ (6)
$G(\sigma,\rho):=\int_{\mathbb{R}^{3}}V_{\sigma}\rho\,dx=\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\rho(x)\sigma(y)}{|x-y|}\,dxdy$ (7)
Here $U(\rho)$ is the internal energy, and $G(\rho,\rho)$ is the gravitational potential energy (also called gravitational interaction energy).
@@
$E(\rho)=E_{0}(\rho)+\frac{1}{2}\int_{\mathbb{R}^{3}}\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)\,dx$ (8)
@@
In this paper we mainly consider the non-rotating case ($L=0$ or $J=0$, reflected in the subscript 0{0} in $E_{0}(\rho)$). This is not only for the sake of simplicity but also helps to produce more results such as uniqueness result in subsection 2.3.
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [17, Section 2]. Hence $E_{0}(\rho)$ is well-defined.
@@
$mR\left(\mathbb{R}^{3}\right)$ $=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=m\right\}$ (11)
@@
From Remark 2.4, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, selecting the specific exponent $\frac{4}{3}$ can also be shown as a natural assumption if we hope the minimal energy $\inf\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)$ to be finite. We will discuss this later (Proposition 2.16, see also [4, Section 1,6 and 8], and [46, Section 2]). We will also show actually $\frac{4}{3}$ is the critical number such that the maximal energy is finite (Corollary 2.18).
@@
Since the energy is translation-invariant, we may just consider looking for an energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, where
$\bar{x}(\rho):=\frac{\int_{\mathbb{R}^{3}}x\rho(x)\,dx}{\int_{\mathbb{R}^{3}}\rho(x)\,dx}$ (13)
@@ 2.2 Existence Results for Non-rotating Bodies
Recall that in the calculus of variations, if an energy functional has a minimizer, then this minimizer must satisfy the Euler–Lagrange equation [^1] When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, [17, Section 2, Appendix A].. Moreover, it turns out that the Euler–Lagrange equation is equivalent, in a certain sense, to the Euler–Poisson equation (see, for example, [17, Section 2] or [33]). Based on this observation, we have Auchmuty, Beals, Lieb and Yau’s results [4, 42]. They not only show the existence of one-body and non-rotating solutions of Euler-Poisson equations, but also indicate some useful properties. McCann lists those results as the following theorem:
@@
    on all of $\mathbb{R}^{3}$ and a single Lagrange multiplier $\lambda_{m}\left\{\begin{array}[]{l}<0,m>0\\ =0,m=0\end{array}\right$.. Here $\cdot$$_{+}$ is the nonnegative (positive) part function defined as $\lambda]_{+}:=\max\{\lambda,0\}$;
@@
Remark 2.8.
Notice Auchmuty and Beals only concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A], whereas in Theorem 2.6 (ix) we establish a stronger result—namely, that the equation (EP’) holds over the entire $\mathbb{R}^{3}$ space.
Given that $\rho$ is an energy minimizer, to exploit this property better, we introduce the following perturbation set $P_{\infty}(\rho)$(the motivation behind its definition and related properties can be found in [17, Section 4]), which depends on $\rho$:
@@
Lemma 2.9 (Differentiability of Energy $E_{0}(\rho)$ [17, Section 5]).
Given $\rho\in m{R}^{3}$ with $U(\rho)<\infty$, we have $E_{0}(\rho)$ is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. The derivative at $\rho$ is $E_{0}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{0}^{\prime}(\rho)(\sigma)=\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\sigma\,dx$ [^2] To remain consistent with the notation in [4, Section 4], we use $E_{0}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{0}^{\prime}(\rho)$ is given by
@@
To apply Lemma 2.9, we need to verify $U(\rho)<\infty$. It turns out if $\rho\in L^{1}\cap L^{\infty}$, or $\rho$ minimizes $E_{0}(\rho)$ locally (under the topology induced by Wasserstein $L^{\infty}$ distance) or globally on $m{R}\left(\mathbb{R}^{3}\right)$, then $U(\rho)<\infty$, see [17, Section 5].
@@
To prove the existence of a minimizer  [^3] It is natural to consider direct method [19]. However, so far, we only know $mR\left(\mathbb{R}^{3}\right)$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, which only implies weak star convergence. It is not enough to show the weak continuity of $G(\rho,\rho)$ and then the weak lower semi-continuity of $E_{0}(\rho)$ with respect to $\rho$ on $W_{\infty}\coloneq m{R}\left(\mathbb{R}^{3}\right)$, see Proposition A.5 or Proposition A.6. Hence, the direct method may not be valid on $W_{\infty}$. We need to construct a constrained admissible class $W_{R}$., we first construct a constrained admissible class:
@@
Here $B_{R}(\cdot)$ is the open ball defined by $B_{R}(\cdot):=\left\{x\in\mathbb{R}^{3}\mid|x-\cdot|<R\right\}$. The axial symmetry and centered condition in (16) ensure that the center of mass of $\rho_{R}$ is $(0,0,0)^{T}$. Notice that $B_{R}(0)$ is a bounded set, by Proposition A.5 or Proposition A.6, we can show the weak continuity of $G(\rho,\rho)$ with respect to $\rho$ on $W_{R}$. Then one can apply direct method to show the existence of a minimizer $\sigma_{R}$ (depending on $m$) on $W_{R}$. In order to show such minimizer $\sigma_{R}$ is truly a minimizer on the whole admissible class $W_{\infty}$ for sufficiently large $R$, we need the following 4 lemmas.
@@
    Show $\forall R>\widetilde{R}_{0}$, $E_{0}\left(\sigma_{R}\right)\leq\alpha<0$, $\lambda_{R}\leq l<0$, where $\lambda_{R}$ is the corresponding Lagrange multiplier. (Lemma 2.24)
@@
The proofs of the above four lemmas are postponed to later text. Once these lemmas are established, we obtain a minimizer which satisfies Parts (iv) (v). Therefore, Part (i) holds true with finite $e_{0}(m)$ by the second lemma. With the strategies similar to that employed in the proofs of the above lemmas, one can show Part (iv) and Part (v) actually hold true for any minimizer $\sigma_{m}$ (observe that the last term in (25) can be replaced by $\frac{m}{4R}$ since $\int_{B_{R}(0)}\sigma_{m}\,dx\geq\frac{m}{2}$ for sufficiently large $R$). More precisely, the proof of Part (v) also relies on Part (vi) to show that outside the ball of radius $R_{0}(m)$, $\sigma_{m}$ is not only almost everywhere zero but truly zero everywhere. Moreover, we have the fact that the Lagrange multiplier $\lambda_{m}$ is negative, which follows from the similar arguments in the proof of Step (1) (Lemma 2.21) showing $\lambda_{R}<0$, see also Remark 2.25.
Now we have the existence of the minimizer $\sigma_{m}$. Since $\sigma_{m}$ is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, the set of admissible perturbations here is exactly $P_{\infty}\left(\sigma_{m}\right)$. One can apply (generalized) Lagrange multiplier theorem ([4, Section 2] or [17, Appendix A]), together with the fact that there are sufficiently many functions in $P_{\infty}(\sigma_{m})$, to obtain the following relations (see also [4, Lemma 2] or [17, Section 5]):
@@
$E_{0}^{\prime}\left(\sigma_{m}\right)=\lambda_{m}\text{ a.e. on }\{\sigma_{m}(x)>0\}.$ (18)
@@
$A^{\prime}\left(\sigma_{m}(x)\right)\leq V_{\sigma_{m}}(x)+\lambda_{m}\text{ a.e. on }\{\sigma_{m}(x)>0\}$ (19)
Due to Lemma 2.2 (iii), notice $\lambda_{m}<0$, we can obtain $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. As in [17, Section 2], we can show the bound of $\|\sigma_{m}\|_{L^{p}}$ and $\left\|V_{\sigma_{m}}\right\|_{L^{\widetilde{p}}}$ with $p$ and $\widetilde{p}$ increasing alternately and reaching infinity in finite steps, and then $V_{\sigma_{m}}$ is continuously differentiable (bootstrap method). On the other hand (17)(18) implies
@@
where $\phi=\left(A^{\prime}\right)^{-1}$ is continuous. This means $\sigma_{m}$ has a continuous representative. Moreover, if $P(\rho)$ satisfies (F4), then $\phi\in C^{1}((0,\infty))$ by Lemma 2.2 (vii), which implies $\sigma_{m}\in C^{1}\left(\left\{\sigma_{m}>0\right\}\right)$ as shown in [17, Section 2]. (Part (vi)).
@@
Part (iii) comes from a strong rearrangement inequality in Lieb [43, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [54]. One can also refer to [46, Proof of Theorem 6.1] or [16, Section 6]. (Part (iii))